OpenStudy (swissgirl):

Transalate the following english statement into symbolic sentences with Quantifiers. Between any integer and any larger integer, there is a real number. The universe is Real numbers

OpenStudy (kinggeorge):

\[\forall n,k\in\mathbb{Z} \;\;\;\exists x\in\mathbb{R} \;\;\;n<x<n+k\]

OpenStudy (anonymous):

\[\forall a\in\mathbb{Z}\forall b\in\mathbb{Z}\exists c\in\mathbb{R} a<b\implies a<c<b\]

OpenStudy (anonymous):

Yeah, either of those two will work.

OpenStudy (swissgirl):

yuppp ok ill medal nbouscal and nbouscal will medal king

OpenStudy (kinggeorge):

Actually mine is a bit wrong. One second.

OpenStudy (anonymous):

Oh, yeah it is. k could be negative.

OpenStudy (kinggeorge):

\[\forall n\in\mathbb{Z}\;\;\;\forall k\in\mathbb{N} \;\;\;\exists x\in\mathbb{R} \;\;\;n<x<n+k\]

OpenStudy (swissgirl):

Thanks GUYS :)))))))